# Gold after Randomized Sand: Model-X Split Knockoffs for Controlled Transformation Selection

### v1

@@ Abstract
Controlling the False Discovery Rate (FDR) in variable selection is crucial for reproducibility and preventing over-selection, particularly with the increasing prevalence of predictive modeling. The Split Knockoff method, a recent extension of the canonical Knockoffs framework, offers finite-sample FDR control for selecting sparse transformations, finding applications across signal processing, economics, information technology, and the life sciences. However, its current formulation is limited to fixed design settings, restricting its use to linear models. The question of whether it can be generalized to random designs, thereby accommodating a broader range of models beyond the linear case – similar to the Model-X Knockoff framework – remains unanswered. A major challenge in addressing transformational sparsity within random design settings lies in reconciling the combination of a random design with a deterministic transformation. To overcome this limitation, we propose the Model-X Split Knockoff method. Our method achieves FDR control for transformation selection in random designs, bridging the gap between existing approaches. This is accomplished by introducing an auxiliary randomized design that interacts with both the existing random design and the deterministic transformation, enabling the construction of Model-X Split Knockoffs. Like the classical Model-X framework, our method provides provable finite-sample FDR control under known or accurately estimated covariate distributions, regardless of the conditional distribution of the response. Importantly, it guarantees at least the same selection power as Model-X Knockoffs when both are applicable. Empirical studies, including simulations and real-world applications to Alzheimer’s disease imaging and university ranking analysis, demonstrate robust FDR control and suggest improved selection power over the original Model-X approach.
@@ 1 Introduction
In another direction, Candes et al. (2018) introduced the Model-X Knockoff framework, which enables FDR control in random design settings. This method leverages knowledge of the marginal distribution of the design matrix and identifies the Markov blanket – the smallest subset of variables that renders the outcome independent of all others – allowing variable selection in a wide range of statistical models. The Model-X framework represents a significant advance, accommodating diverse modeling scenarios and inspiring substantial research (Barber et al., 2020, Romano et al., 2020, Ren and Candès, 2023, Ren and Barber, 2024).
@@
To address this gap, this paper introduces the Model-X Split Knockoff method, which extends the Model-X Knockoff framework to handle transformational sparsity in random design settings. This novel method provides FDR control with finite-sample guarantees under transformational sparsity across a broad range of statistical models, beyond the linear settings considered in earlier works. Moreover, as will be demonstrated, the Model-X Split Knockoff method achieves at least the same selection power as the original Model-X Knockoff method and often outperforms it in empirical applications when both methods are applicable.
@@
where $f:\mathbb{R}\to\mathbb{R}$ is the link function, and $D\in\mathbb{R}^{m\times p}$ is a predetermined linear transformation matrix that induces sparsity in the transformed coefficient vector $\gamma^{*}\in\mathbb{R}^{m}$. Transformation selection focuses on identifying the support set of $\gamma^{*}$, i.e. $\mathcal{H}_{1}^{\gamma}=\{i:\gamma^{*}_{i}\neq 0\}$, rather than directly identifying the support of $\beta^{*}\in\mathbb{R}^{p}$. For example, in multiple comparison problems, $D$ is the graph difference matrix such that $(D\beta^{*})(i,j):=\beta^{*}_{i}-\beta^{*}_{j}$ for each pair $(i,j)$ in an edge set. In this case, $\gamma^{*}_{i}=0$ corresponds to the null hypothesis $\mathcal{H}^{i,j}:\beta_{i}^{*}=\beta_{j}^{*}$, meaning that covariate $X_{i}$ has the same conditional effect as $X_{j}$.
@@
Since the proposed method is grounded in the Model-X framework and does not require specifying the conditional distribution $P_{Y|X}$, it accommodates more complex models beyond simple linear settings. Additionally, this separation provides greater flexibility to optimize selection power compared to the original Model-X Knockoff method when $D=I$, as it imposes no assumptions on the distribution of $A$. Benefiting from this flexibility, the Model-X Split Knockoff method is guaranteed to achieve at least the same power as the Model-X Knockoff method, and empirical results demonstrate that it often achieves superior selection power in practice.
@@
To evaluate the performance of our proposed method, we conduct extensive experiments on both simulated and real-world datasets, including data for Alzheimer’s Disease and world college rankings. Simulation experiments using both regression and pairwise comparison data from random graphs validate the method’s effectiveness. For the Alzheimer’s Disease dataset, the method successfully identifies abnormal regions and connections associated with the disease. Furthermore, it identifies important and significant gaps between college pairs in the world college ranking dataset.
The paper is organized as follows: Section 2 introduces the methodology of Model-X Split Knockoff, including specific construction methods for canonical examples. The theoretical results for FDR control of the proposed method are presented in Section 3. Simulation experiments are detailed in Section 4, while experiments on real-world datasets are discussed in Section 5 and Section 6.
@@ 2 Methodology
In this paper, we address the transformational sparsity problem, modeled in Equation (1) and quoted here for convenience:
$\mathbb{E}(Y|X)=f(X\beta^{*}),\quad\gamma^{*}=D\beta^{*}.$
The primary goal of this paper is to estimate the support set of $\gamma^{*}$, defined as $\mathcal{H}^{\gamma}_{1}=\{i:\gamma^{*}_{i}\neq 0\}$, while controlling the false discovery rate. A key challenge arises from the interplay of a random design $X$ and a deterministic transformation $D$, where neither directly maps $\gamma^{*}$ to the outcome $Y$.
@@ 2.1 Variable Splitting and Auxiliary Randomized Design
Recall that the equation holds because $\widetilde{X}^{A}:=X-AD$ and $A\gamma^{*}=D\beta^{*}$.
In this way, we successfully inserted an auxiliary $A$ for the parameter of interest $\gamma^{*}$, and lifted the model (1) into a higher dimensional space including both $\beta$ and $\gamma$. With Equation (3), a Split Knockoff copy $\widetilde{A}$ following certain necessary constraints could be constructed to enable variable selection with the false discovery rate control on $\gamma^{*}$. After such construction, standard procedures of Split Knockoffs as described in Cao et al. (2024a) could be applied to finalize the procedure. In the following section, we will proceed to the step of constructing the Split Knockoff copy.
@@ 2.2 Construction of Model-X Split Knockoffs
Similarly to (Candes et al., 2018, Cao et al., 2024a), the Model-X Split Knockoff requires the construction of so-called Split Knockoff copies to achieve the false discovery rate control which will be shown in Section 3. The random vector $\widetilde{A}=(\widetilde{A}_{1},\dots,\widetilde{A}_{m})$ is called a Split Knockoff copy with respect to $(\widetilde{X}^{A},A)$ if the following conditions are satisfied:
@@
    $(\widetilde{X}^{A},A,\widetilde{A})_{\text{swap}(S)}\overset{\mathcal{D}}{=}(\widetilde{X}^{A},A,\widetilde{A}).$ (4)
@@
Here, $(\widetilde{X}^{A},A,\widetilde{A})_{\text{swap}(S)}$ means swap features in $S$ between $A$ and $\widetilde{A}$. For example, if $A\in R^{2},S=\{1\}$, then
$(\widetilde{X}^{A},A_{1},A_{2},\widetilde{A}_{1},\widetilde{A}_{2})_{\text{swap}(S)}=(\widetilde{X}^{A},\widetilde{A}_{1},A_{2},A_{1},\widetilde{A}_{2}).$
In the cases where the distribution of $X$ is well known, for any choice of $A$, the $\widetilde{A}$ satisfying the exchangeability condition (4) can be constructed in a sequential way as shown in Algorithm 1, similar to Candes et al. (2018).
@@
Generate random variable $A$ and the respective random variable $\widetilde{X}^{A}$
@@
The proof of Algorithm 1 satisfies the exchangeability is given in Supplementary Material Section A.
Although Algorithm 1 gives the general construction of the Split Knockoff copy $\widetilde{A}$, it might be facing computational challenges in practice as the computational cost enlarges quickly as $m$ enlarges. Moreover, the generated Split Knockoff copies can be highly correlated which harms the selection power. In the following, we deal with two canonical examples to present refined constructions of Split Knockoff copies, which leads to lower computational costs and better selection power.
@@ 2.2.1 Construction for Normal Distributions
In the cases where the distribution of $X$ follows or can be well-approximated by the normal distribution $\mathcal{N}(0,\Sigma_{X})$, the nature choice of the pair $(A,\widetilde{A})$ is to generate it from a joint normal distribution of $(X,A,\widetilde{A})$, as suggested by Candes et al. (2018). Specially, as we shall shown later, among all choices of joint normal distributions, the choices which makes the “intercept” $\widetilde{X}^{A}$ independent from the auxiliary design $A$ are expected to achieve higher selection power.
The generation process begins with $A$ being drawn from
$(X,A)\sim\mathcal{N}\left\{\left(\begin{array}[]{c}\mathbf{0}_{p}\\ \mathbf{0}_{m}\end{array}\right),\left(\begin{array}[]{cc}\Sigma_{X}&\Sigma_{AX}^{T}\\ \Sigma_{AX}&\Sigma_{A}\end{array}\right)\right\},$ (9)
where $\Sigma_{A}$ and $\Sigma_{AX}$ are two suitable matrices which could be used to optimize the selection power. Then by definition of $\widetilde{X}^{A}=X-AD$, there holds
$(\widetilde{X}^{A},A)\sim\mathcal{N}\left\{\left(\begin{array}[]{c}\mathbf{0}_{p}\\ \mathbf{0}_{m}\end{array}\right),\left(\begin{array}[]{cc}\Sigma_{\widetilde{X}^{A}}&\Sigma_{AX}^{T}-D^{T}\Sigma_{A}\\ \Sigma_{AX}-\Sigma_{A}D&\Sigma_{A}\end{array}\right)\right\}.$ (14)
where $\Sigma_{\widetilde{X}^{A}}=\Sigma_{X}-\Sigma_{XA}^{T}D-\Sigma_{XA}D^{T}+D^{T}\Sigma_{A}D$. Then the conditional distribution of $A|\widetilde{X}^{A}$ evaluated at ${\widetilde{X}^{A}}=\widetilde{x}^{A}$ in Algorithm 1 is
$A|(\widetilde{X}^{A}=\widetilde{x}^{A})\sim\mathcal{N}\left\{(\Sigma_{AX}-\Sigma_{A}D)\Sigma_{\widetilde{X}^{A}}^{-1}(\widetilde{x}^{A})^{T},\Sigma_{A}-(\Sigma_{AX}-\Sigma_{A}D)\Sigma_{\widetilde{X}^{A}}^{-1}(\Sigma_{AX}^{T}-D^{T}\Sigma_{A})\right\}$ (15)
Then conditional on the observation $\widetilde{X}^{A}=\widetilde{x}^{A}$, the Split Knockoff copy $\widetilde{A}$ could be generated from the joint distribution
$(A,\widetilde{A})|(\widetilde{X}^{A}=\widetilde{x}^{A})\sim\mathcal{N}\left\{\left(\begin{array}[]{c}(\Sigma_{AX}-\Sigma_{A}D)\Sigma_{\widetilde{X}^{A}}^{-1}(\widetilde{x}^{A})^{T}\\ (\Sigma_{AX}-\Sigma_{A}D)\Sigma_{\widetilde{X}^{A}}^{-1}(\widetilde{x}^{A})^{T}\end{array}\right),\left(\begin{array}[]{cc}\widetilde{\Sigma}_{A}&\widetilde{\Sigma}_{A}-\mathrm{diag}(\mathbf{s})\\ \widetilde{\Sigma}_{A}-\mathrm{diag}(\mathbf{s})&\widetilde{\Sigma}_{A}\end{array}\right)\right\},$ (20)
where $\widetilde{\Sigma}_{A}=\Sigma_{A}-(\Sigma_{AX}-\Sigma_{A}D)\Sigma_{\widetilde{X}^{A}}^{-1}(\Sigma_{AX}^{T}-D^{T}\Sigma_{A})$, $s\in\mathbb{R}^{m}$ is an non-negative vector which makes the above covariance matrix semi-positive definite. The proof of the argument that this construction satisfies the exchangeability can be directly obtained from the normal distributions and will be omitted.
Now we are going to specify the choice of $\Sigma_{A}$ and $\Sigma_{AX}$ in order to optimize the selection power.
@@
    Choice of $\Sigma_{AX}$. As discussed in Barber and Candès (2015), Candes et al. (2018), the non-negative vector $s$ in Equation (20) should be as large as possible to maximize the selection power. It could be verified that $s$ needs to satisfy
    $s\preceq 2\widetilde{\Sigma}_{A}\preceq 2\Sigma_{A},$
    where for the second inequality, the equation holds if and only if $\Sigma_{AX}=\Sigma_{A}D$. This gives us evidence that taking $\Sigma_{AX}=\Sigma_{A}D$, i.e. $\widetilde{X}^{A}$ independent from $A$ in (14), helps optimize the selection power.
@@
In summary, the choice of $\Sigma_{A}$ and $\Sigma_{AX}$ to optimize the selection power is $\Sigma_{A}=\alpha I_{m}$ and $\Sigma_{AX}=\alpha D$. In this case, the joint distribution over $(X,A)$ is
$(X,A)\sim\mathcal{N}\left\{\left(\begin{array}[]{c}\mathbf{0}_{p}\\ \mathbf{0}_{m}\end{array}\right),\left(\begin{array}[]{cc}\Sigma_{X}&\alpha D^{\top}\\ \alpha D&\alpha I_{m}\end{array}\right)\right\}.$ (25)
where $\alpha$ satisfies $\Sigma_{X}-\alpha D^{T}D\succeq 0$. In this setting, $\widetilde{A}$ is generated from
$(A,\widetilde{A})\sim\mathcal{N}\left\{\left(\begin{array}[]{c}\mathbf{0}_{m}\\ \mathbf{0}_{m}\end{array}\right),\left(\begin{array}[]{cc}\alpha I_{m}&\alpha I_{m}-\mathrm{diag}(\mathbf{s})\\ \alpha I_{m}-\mathrm{diag}(\mathbf{s})&\alpha I_{m}\end{array}\right)\right\}.$ (30)
This specific choice of $A$ and $\widetilde{A}$ generates independent designs for features of interests, which enhances the selection power as validated by experiments in Section 4.1. Moreover, when the covariance matrix $\Sigma_{X}$ is unknown, one can see a good estimation of such covariance matrices leads to a robust FDR control in Section 3.2.

### v2

@@ Abstract
Controlling the False Discovery Rate (FDR) is critical for reproducible variable selection, especially given the prevalence of complex predictive modeling. The recent Split Knockoff method, an extension of the canonical Knockoffs framework, offers finite-sample FDR control for selecting sparse transformations but is limited to linear models with fixed designs. Extending this framework to random designs, which would accommodate a much broader range of models, is challenged by the fundamental difficulty of reconciling a random covariate design with a deterministic linear transformation. To bridge this gap, we introduce Model-X Split Knockoffs. Our method achieves robust FDR control for transformation selection in random designs by introducing a novel auxiliary randomized design. This key innovation effectively mediates the interaction between the random design and the deterministic transformation, enabling the construction of valid knockoffs. Like the classical Model-X framework, our approach provides provable, finite-sample FDR control under known or accurately estimated covariate distributions, regardless of the response’s conditional distribution. Importantly, it guarantees at least the same, and often superior, selection power as standard Model-X Knockoffs when both are applicable. Empirical studies, including simulations and real-world applications to Alzheimer’s disease imaging and university ranking analysis, demonstrate robust FDR control and improved statistical power.
@@ 1 Introduction
In another direction, Candès et al. (2018) introduced the Model-X Knockoff framework, which enables FDR control in random design settings. This method leverages knowledge of the marginal distribution of the design matrix and identifies the Markov blanket – the smallest subset of variables that renders the outcome independent of all others – allowing variable selection in a wide range of statistical models. The Model-X framework represents a significant advance, accommodating diverse modeling scenarios and inspiring substantial research (Barber et al., 2020, Romano et al., 2020, Ren and Candès, 2023, Ren and Barber, 2024).
@@
To address this gap, this paper introduces the Model-X Split Knockoff method, which extends the Model-X Knockoff framework to handle transformational sparsity in random design settings. This novel method provides FDR control with finite-sample guarantees under transformational sparsity across a broad range of statistical models, beyond the linear settings considered in earlier works (Cao et al., 2024a, b). Moreover, as will be demonstrated, the Model-X Split Knockoff method achieves at least the same selection power as the original Model-X Knockoff method and often outperforms it in empirical applications when both methods are applicable.
@@
where $f:\mathbb{R}\to\mathbb{R}$ is the link function, and $D\in\mathbb{R}^{m\times p}$ is a predetermined linear transformation matrix that induces sparsity in the transformed coefficient vector $\gamma^{*}\in\mathbb{R}^{m}$. Transformation selection focuses on identifying the support set of $\gamma^{*}$, i.e. $\mathcal{H}_{1}^{\gamma}=\{i:\gamma^{*}_{i}\neq 0\}$, rather than directly identifying the support of $\beta^{*}\in\mathbb{R}^{p}$. For example, in multiple comparison problems, $D$ is the graph difference matrix such that $(D\beta^{*})_{(j,k)}:=\beta^{*}_{j}-\beta^{*}_{k}$ for each pair $(j,k)$ in an edge set. In this case, $\gamma^{*}_{(j,k)}=0$ corresponds to the null hypothesis $\mathcal{H}^{j,k}_{0}:\beta_{j}^{*}=\beta_{k}^{*}$, meaning that covariate $X_{j}$ has the same conditional effect as $X_{k}$.
@@
Since the proposed method is grounded in the Model-X framework and does not require specifying the conditional distribution $P_{Y|X}$, it accommodates more complex models beyond simple linear ones. Additionally, this separation provides greater flexibility to optimize selection power compared to the original Model-X Knockoff method when $D=I$, as it imposes no assumptions on the distribution of $A$. Benefiting from this flexibility, the Model-X Split Knockoff method is guaranteed to achieve at least the same power as the Model-X Knockoff method, and empirical results demonstrate that it often achieves superior selection power in practice.
@@
To evaluate the performance of our proposed method, we conduct experiments on both simulated and real-world datasets, including the Alzheimer’s Disease and world college rankings. Simulation experiments using both regression and pairwise comparison data from random graphs validate the method’s effectiveness. For the Alzheimer’s Disease dataset, the method successfully identifies abnormal regions and connections associated with the disease. Furthermore, it identifies important and significant gaps between college pairs in the world college ranking dataset.
We introduce the Model-X Split Knockoff methodology in Section 2 and its theoretical results for FDR control in Section 3. Simulation experiments are detailed in Section 4, followed by applications on real-world datasets in Section 5 and Section 6.
@@ 2 Methodology
In this paper, we address the transformational sparsity problem, modeled in Equation (1) and quoted here for convenience: $\mathbb{E}(Y|X)=f(X\beta^{*})$ and $\gamma^{*}=D\beta^{*}$. The primary goal of this paper is to estimate the support set of $\gamma^{*}$, defined as $\mathcal{H}^{\gamma}_{1}=\{i:\gamma^{*}_{i}\neq 0\}$, while controlling the false discovery rate. A key challenge arises from the interplay of a random design $X$ and a deterministic transformation $D$, where neither directly maps $\gamma^{*}$ to the outcome $Y$.
@@ 2.1 Variable Splitting and Auxiliary Randomized Design
Recall that the equation holds because $\widetilde{X}^{A}:=X-AD$ and $A\gamma^{*}=AD\beta^{*}$.
In this way, we successfully inserted an auxiliary $A$ for the parameter of interest $\gamma^{*}$, and lifted the model (1) into a higher-dimensional space including both $\beta^{*}$ and $\gamma^{*}$. With Equation (3), a split knockoff copy $\widetilde{A}$ could be constructed following certain necessary constraints to allow variable selection with the control of the false discovery rate in $\gamma^{*}$. After such a construction, standard Split Knockoff procedures as described in Cao et al. (2024a) could be applied to finalize the procedure. In the following section, we will proceed to the step of constructing the Split Knockoff copy.
@@ 2.2 Construction of Model-X Split Knockoffs
Similarly to (Candès et al., 2018, Cao et al., 2024a), the Model-X Split Knockoff requires the construction of Split Knockoff copies to achieve the false discovery rate control to be shown in Section 3. The random [^1] To account for the randomness introduced by our construction, one can apply the e-value scheme of (Wang and Ramdas, 2022, Lee and Ren, 2024) detailed in Supplementary Material Section H. vector $\widetilde{A}=(\widetilde{A}_{1},\dots,\widetilde{A}_{m})$ is called a Split Knockoff copy with respect to $(\widetilde{X}^{A},A)$ if the following conditions hold.
@@
[…]
@@
Here, $(\widetilde{X}^{A},A,\widetilde{A})_{\text{swap}(S)}$ means to swap the features in $S$ between $A$ and $\widetilde{A}$. For example, if $A\in\mathbb{R}^{2},S=\{1\}$, then $(\widetilde{X}^{A},A_{1},A_{2},\widetilde{A}_{1},\widetilde{A}_{2})_{\text{swap}(S)}=(\widetilde{X}^{A},\widetilde{A}_{1},A_{2},A_{1},\widetilde{A}_{2})$.
In the cases where the distribution of $X$ is well known, for any choice of $A$, $\widetilde{A}$ can be constructed sequentially to satisfy the exchangeability condition (4), as shown in Algorithm 1, which is similar to Candès et al. (2018). The proof of Algorithm 1 satisfies the exchangeability is given in Supplementary Material Section A.
@@
Generate random variable AA and the respective random variable $\widetilde{X}^{A}$
@@
Although Algorithm 1 gives the general construction of the Split Knockoff copy $\widetilde{A}$, it might face computational challenges in practice as the computational cost increases rapidly as $m$ increases. Moreover, the generated Split Knockoff copies can be highly correlated, which harms the selection power. In the following, we deal with two canonical examples to present refined constructions of Split Knockoff copies, which leads to lower computational costs and better selection power.
@@ 2.2.1 Construction for Normal Distributions
[…]
The generation process begins with $A$ drawn from $(X,A)\sim\mathcal{N}(\mu,\Sigma)$, where $\mu=\mathbf{0}_{m+p}$ and $\Sigma=(\Sigma_{X}\;\;\Sigma_{AX}^{\top};\;\;\Sigma_{AX}\;\;\Sigma_{A})$, with $\Sigma_{A}$ and $\Sigma_{AX}$ being two suitable matrices which could be used to optimize the selection power. With some calculations detailed in Supplementary Material Section F, conditional on the observation $\widetilde{X}^{A}=\widetilde{x}^{A}$, the Split Knockoff copy $\widetilde{A}$ could be generated from the joint distribution
$(A,\widetilde{A})|(\widetilde{X}^{A}=\widetilde{x}^{A})\sim\mathcal{N}\left\{\left(\begin{array}[]{c}\widetilde{\mu}_{\widetilde{x}^{A}}\\ \widetilde{\mu}_{\widetilde{x}^{A}}\end{array}\right),\left(\begin{array}[]{cc}\widetilde{\Sigma}_{A}&\widetilde{\Sigma}_{A}-\mathrm{diag}(\mathbf{s})\\ \widetilde{\Sigma}_{A}-\mathrm{diag}(\mathbf{s})&\widetilde{\Sigma}_{A}\end{array}\right)\right\},$ (9)
[…]
We are going to specify $\Sigma_{A}$ and $\Sigma_{AX}$ in order to optimize the selection power.
@@
    Choice of $\Sigma_{AX}$. As discussed in (Barber and Candès, 2015, Candès et al., 2018), the vector $\mathbf{s}$ in Equation (9) should be as large positive as possible to maximize the selection power. It could be verified that $\mathbf{s}$ needs to satisfy $\mathrm{diag}(\mathbf{s})\preceq 2\widetilde{\Sigma}_{A}\preceq 2\Sigma_{A},$ where the equality of the second inequality occurs if and only if $\Sigma_{AX}=\Sigma_{A}D$. This gives us evidence that taking $\Sigma_{AX}=\Sigma_{A}D$, i.e. $\widetilde{X}^{A}\perp A$ as $cov(\widetilde{X}^{A},A)=\Sigma_{AX}-\Sigma_{A}D=0$, helps improve the selection power.
@@
[…]
$\mathcal{N}\left\{\left(\begin{array}[]{c}\mathbf{0}_{p}\\ \mathbf{0}_{m}\\ \mathbf{0}_{m}\end{array}\right),\left(\begin{array}[]{ccc}\Sigma_{X}&\alpha D^{\top}&\alpha D^{\top}-\mathrm{diag}(\mathbf{s})D^{\top}\\ \alpha D&\alpha I_{m}&\alpha I_{m}-\mathrm{diag}(\mathbf{s})\\ \alpha D-\mathrm{diag}(\mathbf{s})D&\alpha I_{m}-\mathrm{diag}(\mathbf{s})&\alpha I_{m}\end{array}\right)\right\}.$ (16)
This choice of $A$ and $\widetilde{A}$ ensures $(A,\widetilde{A})\perp\widetilde{X}^{A}$, which enhances the selection power as validated by experiments in Section 4.1. Moreover, when $\Sigma_{X}$ is unknown, one can see a good estimation of $\Sigma_{X}$ leads to a robust FDR control in Section 3.